6.) Kyle is driving his car at 15 m/s when he noticed that he is approaching a stalled truck that is approximately 35 m away. He steps on the brakes and the car stop 4.0 m away from the stalled truck. At what rate did the car slow down? Assume that Kyle's reaction time is 2.0 s. Reaction time is the time for a person to notice, think, and act in response to a given situation.
1
Expert's answer
2019-08-22T09:27:44-0400

Distance traveled by car before any reaction=

"u \times t=15 \times 2=30 m"

Distance remaining between car and truck=5 m

Car stops 4 m away from truck,so distance traveled by car is 1 m and velocity of car comes down to zero.

By equation of motion,

"v^2-u^2=2as.......(v=0)"

"a=-\frac{u^2}{2s}=-\frac{15^2}{2 \times 1}=-112.5 m/sec^2"

("- ve" sign shows retardation)